OpenStudy (anonymous):

Find the slope of a line that passes through the two points (4,-3) and (-8,0).

OpenStudy (anonymous):

\[m=\frac{\Delta Y}{\Delta X}\]\[m=\frac{-3-0}{4-(-8)}\]The slope is:\[m=\frac{3}{4}\]

OpenStudy (anonymous):

Thank-you!, can you help me with another? :x

OpenStudy (anonymous):

@D3xt3R

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Write an equation of a horizontal line passing through the two points (-1,5) and (0,5). Do not use any spaces in your answer. @D3xt3R

OpenStudy (anonymous):

This horizontal line is an perpendicular line, so:\[y-y_o=m*(x-x_o)\]\[m_1=\frac{5-5}{-1-0}\]\[m_1=0\]The horizontal line should have an slope \[m_2=-\frac{1}{m_1}\]\[m_2=\frac{1}{0}=\nexists\]But you can write as a constant \[x=Any~~~number\]